# Faster and Stronger: When ANN-SNN Conversion Meets Parallel Spiking Calculation

### v1

@@ Abstract
Spiking Neural Network (SNN), as a brain-inspired and energy-efficient network, is currently facing the pivotal challenge of exploring a suitable and efficient learning framework. The predominant training methodologies, namely Spatial-Temporal Back-propagation (STBP) and ANN-SNN Conversion, are encumbered by substantial training overhead or pronounced inference latency, which impedes the advancement of SNNs in scaling to larger networks and navigating intricate application domains. In this work, we propose a novel parallel conversion learning framework, which establishes a mathematical mapping relationship between each time-step of the parallel spiking neurons and the cumulative spike firing rate. We theoretically validate the lossless and sorting properties of the conversion process, as well as pointing out the optimal shifting distance for each step. Furthermore, by integrating the above framework with the distribution-aware error calibration technique, we can achieve efficient conversion towards more general activation functions or training-free circumstance. Extensive experiments have confirmed the significant performance advantages of our method for various conversion cases under ultra-low time latency. To our best knowledge, this is the first work which jointly utilizes parallel spiking calculation and ANN-SNN Conversion, providing a highly promising approach for SNN supervised training.
1 Introduction
Recently, parallel spiking neurons (Fang et al., 2023) is favored by researchers due to its high-speed calculation ability, but current research around it is generally limited to the field of STBP Training. In fact, the training precision of parallel neurons heavily relies on parameter initialization and cannot effectively contribute to the training memory explosion problem of SNNs. Instead, it is likely to play a crucial role in the conversion series methods with higher time latency. In this work, we innovatively combine ANN-SNN Conversion with parallel computing to propose a lossless and high-speed universal parallel conversion framework. The specific contributions are as follows:
@@ 2 Related Works
STBP training for SNNs. As a significant learning algorithm that can achieve relatively superior performance for SNNs within ultra-low time latency, Wu et al. (2018) and Neftci et al. (2019) integrated Back-propagation through Time (BPTT) with surrogate gradient to pioneer the concept of STBP. On this basis, researchers have successively combined STBP algorithm with novel BatchNorm (BN) layers (Zheng et al., 2021; Jiang et al., 2024), residual blocks (Fang et al., 2021; Hu et al., 2024), objective learning functions (Li et al., 2021b; Deng et al., 2022; Guo et al., 2022), multi-dimensional attention mechanisms Qiu et al. (2024), Transformer blocks (Zhou et al., 2023; Shi et al., 2024) and advanced spiking models (Yao et al., 2022; Hao et al., 2024a), thereby extending SNN models to various application scenarios (Ren et al., 2023; Liao et al., 2024; Yao et al., 2024). In addition, to effectively optimize GPU memory overhead and power consumption of STBP training, multiple variant learning frameworks have been further explored, such as time-based learning (Mostafa, 2017) and online learning (Xiao et al., 2022). However, at present, aforementioned schemes still have bottlenecks in learning performance.
@@
ANN-SNN Conversion. Compared to STBP training, ANN-SNN Conversion merely needs to replace the activation function modules of the pretrained ANN model with spiking neurons layer by layer to obtain the converted SNN (Cao et al., 2015; Diehl et al., 2015; Sengupta et al., 2019), which not only economizes training load, but also ensures that the converted SNN has sufficiently superior performance upper-bound. Rueckauer et al. (2017) and Han et al. (2020) realized that the soft-reset mechanism is more suitable for conversion learning algorithms, while Deng and Gu (2021) proposed a shiftable ReLU function to better adapt the distribution of the spike firing rate. Based on the above findings, Bu et al. (2022) proposed an advanced quantization activation function and theoretically validated its ability to losslessly simulate the average firing rate within any time latency from the perspective of mathematical expectation. However, despite achieving better performance than STBP within sufficient time-steps for the same network backbone (Li et al., 2022), converted SNNs still suffer from significant performance degradation under the condition of ultra-low latency (Rathi and Roy, 2021). Therefore, subsequent researchers further proposed more radical error rectification strategies, including setting silent states for specific neurons (Hao et al., 2023a), shifting initial membrane potential (Hao et al., 2023b), as well as introducing burst and signed spikes (Li and Zeng, 2022; Wang et al., 2022).
@@
Training-Free Conversion. How to effectively convert unprocessed ANNs into SNNs with the lowest possible time latency has also become a recent academic focus in the field of conversion learning. Han and Roy (2020) proposed a novel Temporal-Switch-Coding scheme to cut down the time latency and number of addition operations during the SNN inference stage. From the perspectives of calibrating bias and initial membrane potential, Li et al. (2021a) designed two sets of pipelines to further enhance the performance of converted SNNs within dozens of time-steps. Bu et al. (2024) analyzed the upper-bound of the layer-wise conversion error, then pointed out a training-free threshold balancing strategy, which can be applied to various visual tasks on networks with sequential structure.
@@ 4.2 Towards Universal Conversion Error Rectification
  Algorithm 1 Dynamic Calibration Strategy for DA-QCFS Function
@@ 5 Experiments
Consistent with previous conversion learning works, we conduct performance validation on CIFAR (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets by using two types of network backbones, VGG (Simonyan and Zisserman, 2014) and ResNet (He et al., 2016). We selected multiple methods including STBP Training (Li et al., Dspike; Guo et al., RecDis; Yao et al., GLIF; Jiang et al., TAB), ANN-SNN Conversion (Deng and Gu, OPT; Bu et al., QCFS), Conversion Rectification (Li and Zeng, Burst; Wang et al., SNM; Hao et al., SRP; Hao et al., COS), and Training-Free Conversion (Li et al., SNNC; Bu et al., TBC) as comparison targets. The detailed experimental configuration is provided in Appendix.
@@ Appendix A Appendix
(ii) Since QCFS function has the property of $\mathbb{E}\left(\mathbf{r}_{\textsc{QCFS}}^{l,T}-\mathbf{r}_{\textsc{QCFS}}^{l,\tilde{T}}\right)=\mathbf{0}$ when $\psi^{l}=\theta^{l}/2$, as mentioned in [Bu et al., 2022], combining with the conclusion of (i), we can have:
@@ A.2 Detailed Experimental Configuration
For pretrained QCFS ANN models, we use SGD optimizer [Bottou, 2012], the optimization strategy of Cosine Annealing [Loshchilov and Hutter, 2017] and data augmentation techniques [DeVries and Taylor, 2017, Cubuk et al., 2019], the corresponding hyper-parameter settings are: $\texttt{lr}=0.1,\texttt{wd}=5\times 10^{-4}$ for CIFAR-10, $\texttt{lr}=0.02,\texttt{wd}=5\times 10^{-4}$ for CIFAR-100 and $\texttt{lr}=0.1,\texttt{wd}=1\times 10^{-4}$ for ImageNet-1k. The specific network structure is consistent with [Bu et al., 2022]. Regarding the error calibration technique, we utilize the training dataset as the calibration data to iterate for 1 epoch. The learning momentum $\alpha$ mentioned in Algorithm 4.2 is set to $0.99$.
@@ A.3 Comprehensive Analysis based on Inference Speed and Performance
As shown in Fig.S1, we make a comprehensive comparison between parallel and serial inference in terms of speed and learning accuracy. The serial inference performance under the QCFS conversion framework utilizes the accuracies reported in [Bu et al., 2022]. For relatively simple cases (i.e. CIFAR-100 dataset), IF neuron requires approximately 4×4$\times$4 × time latency of parallel conversion to reach the comparable accuracy, which makes our method achieve 1523×15$\sim$ 23$\times$15 ∼ 23 × actual acceleration ratio.

### v2

@@ Abstract
Spiking Neural Network (SNN), as a brain-inspired and energy-efficient network, is currently facing the pivotal challenge of exploring a suitable and efficient learning framework. The predominant training methodologies, namely Spatial-Temporal Back-propagation (STBP) and ANN-SNN Conversion, are encumbered by substantial training overhead or pronounced inference latency, which impedes the advancement of SNNs in scaling to larger networks and navigating intricate application domains. In this work, we propose a novel parallel conversion learning framework, which establishes a mathematical mapping relationship between each time-step of the parallel spiking neurons and the cumulative spike firing rate. We theoretically validate the lossless and sorting properties of the conversion process, as well as pointing out the optimal shifting distance for each step. Furthermore, by integrating the above framework with the distribution-aware error calibration technique, we can achieve efficient conversion towards more general activation functions or training-free circumstance. Extensive experiments have confirmed the significant performance advantages of our method for various conversion cases under ultra-low time latency. To our best knowledge, this is the first work which jointly utilizes parallel spiking calculation and ANN-SNN Conversion, providing a highly promising approach for SNN supervised training. Code is available at https://github.com/hzc1208/Parallel_Conversion.
Machine Learning, ICML
@@
1 Introduction
Recently, parallel spiking neuron (Fang et al., 2023) is favored by researchers due to its high-speed calculation ability, but current research around it is generally limited to the field of STBP Training. In fact, the training precision of parallel neurons heavily relies on parameter initialization and cannot effectively contribute to the training memory explosion problem of SNNs. Instead, it is likely to play a crucial role in the conversion series methods with higher time latency. In this work, we innovatively combine ANN-SNN Conversion with parallel computing to propose a lossless and high-speed universal parallel conversion framework. The specific contributions are as follows:
@@ 2 Related Works
STBP training for SNNs. As a significant learning algorithm that can achieve relatively superior performance for SNNs within ultra-low time latency, Wu et al. (2018) and Neftci et al. (2019) integrated Back-propagation through Time (BPTT) with surrogate gradient to pioneer the concept of STBP. On this basis, researchers have successively combined STBP algorithm with novel BatchNorm (BN) layers (Zheng et al., 2021; Jiang et al., 2024), residual blocks (Fang et al., 2021; Hu et al., 2024), objective learning functions (Li et al., 2021b; Deng et al., 2022; Guo et al., 2022), multi-dimensional attention mechanisms (Qiu et al., 2024), Transformer blocks (Zhou et al., 2023; Shi et al., 2024) and advanced spiking models (Yao et al., 2022; Hao et al., 2024a), thereby extending SNN models to various application scenarios (Ren et al., 2023; Liao et al., 2024; Yao et al., 2024). In addition, to effectively optimize GPU memory overhead and power consumption of STBP training, multiple variant learning frameworks have been further explored, such as time-based learning (Mostafa, 2017) and online learning (Xiao et al., 2022). However, at present, aforementioned schemes still have bottlenecks in learning performance.
@@
ANN-SNN Conversion. Compared to STBP training, ANN-SNN Conversion merely needs to replace the activation function modules of the pretrained ANN model with spiking neurons layer by layer to obtain the converted SNN (Cao et al., 2015; Diehl et al., 2015; Sengupta et al., 2019), which not only economizes training load, but also ensures that the converted SNN has sufficiently superior performance upper-bound. Rueckauer et al. (2017) and Han et al. (2020) realized that the soft-reset mechanism is more suitable for conversion learning algorithms, while Deng & Gu (2021) proposed a shiftable ReLU function to better adapt the distribution of the spike firing rate. Based on the above findings, Bu et al. (2022) proposed an advanced quantization activation function and theoretically validated its ability to losslessly simulate the average firing rate within any time latency from the perspective of mathematical expectation. However, despite achieving better performance than STBP within sufficient time-steps for the same network backbone (Li et al., 2022), converted SNNs still suffer from significant performance degradation under the condition of ultra-low latency (Rathi & Roy, 2021). Therefore, subsequent researchers further proposed more radical error rectification strategies, including setting silent states for specific neurons (Hao et al., 2023a), calibrating forward temporal bias (Wu et al., 2024), shifting initial membrane potential (Hao et al., 2023b), as well as introducing burst and signed spikes (Li & Zeng, 2022; Wang et al., 2022).
@@
Training-Free Conversion. How to effectively convert unprocessed ANNs into SNNs with the lowest possible time latency has also become a recent academic focus in the field of conversion learning. Han & Roy (2020) proposed a novel Temporal-Switch-Coding scheme to cut down the time latency and number of addition operations during the SNN inference stage. From the perspectives of calibrating bias and initial membrane potential, Li et al. (2021a) designed two sets of pipelines to further enhance the performance of converted SNNs within dozens of time-steps. Bu et al. (2024) analyzed the upper-bound of the layer-wise conversion error, then pointed out a training-free threshold balancing strategy, which can be applied to various visual tasks on networks with sequential structure.
@@ 4.2 Towards Universal Conversion Error Rectification
  Algorithm 1 The overall pseudo-code for universal parallel conversion
@@ 5 Experiments
Consistent with previous conversion learning works, we conduct performance validation on CIFAR (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets by using two types of network backbones, VGG (Simonyan & Zisserman, 2014) and ResNet (He et al., 2016). We selected multiple methods including STBP Training (Li et al., 2021b, Dspike; Guo et al., 2022, RecDis; Yao et al., 2022, GLIF; Jiang et al., 2024, TAB), ANN-SNN Conversion (Deng & Gu, 2021, OPT; Bu et al., 2022, QCFS), Conversion Rectification (Li & Zeng, 2022, Burst; Wang et al., 2022, SNM; Hao et al., 2023a, SRP; Hao et al., 2023b, COS), and Training-Free Conversion (Li et al., 2021a, SNNC; Bu et al., 2024, TBC) as comparison targets. The detailed experimental configuration is provided in Appendix.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62422601, U24B20140, and 62088102), Beijing Municipal Science and Technology Program (Z241100004224004), Beijing Nova Program (20230484362, 20240484703), and National Key Laboratory for Multimedia Information Processing.
Impact Statement
This paper presents work whose goal is to advance the field of SNN supervised learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
(ii) Since QCFS function has the property of $\mathbb{E}\left(\mathbf{r}_{\textsc{QCFS}}^{l,T}-\mathbf{r}_{\textsc{QCFS}}^{l,\tilde{T}}\right)=\mathbf{0}$ when $\psi^{l}=\theta^{l}/2$, as mentioned in (Bu et al., 2022), combining with the conclusion of (i), we can have:
@@
Among them, $\mathbf{r}^{l,T}\rightarrow\mathbf{r}_{\textsc{QCFS}}^{l,T}$ maintains lossless conversion under any precondition, while $\mathbf{r}_{\textsc{QCFS}}^{l,T}\rightarrow\mathbf{r}_{\textsc{QCFS}}^{l,\tilde{T}}$ needs to satisfy the assumption mentioned in (Bu et al., 2022) that the input current follows a uniform distribution within multiple sub-intervals. ∎
@@ A.2 Detailed Experimental Configuration
For pretrained QCFS ANN models, we use SGD optimizer (Bottou, 2012), the optimization strategy of Cosine Annealing (Loshchilov & Hutter, 2017) and data augmentation techniques (DeVries & Taylor, 2017; Cubuk et al., 2019), the corresponding hyper-parameter settings are: $\texttt{lr}=0.1,\texttt{wd}=5\times 10^{-4}$ for CIFAR-10, $\texttt{lr}=0.02,\texttt{wd}=5\times 10^{-4}$ for CIFAR-100 and $\texttt{lr}=0.1,\texttt{wd}=1\times 10^{-4}$ for ImageNet-1k. The specific network structure is consistent with (Bu et al., 2022). Regarding the error calibration technique, we utilize the training dataset as the calibration data to iterate for 1 epoch. The learning momentum $\alpha$ mentioned in Algorithm 4.2 is set to $0.99$.
@@ A.3 Comprehensive Analysis based on Inference Speed and Performance
As shown in Fig.S1, we make a comprehensive comparison between parallel and serial inference in terms of speed and learning accuracy. The serial inference performance under the QCFS conversion framework utilizes the accuracies reported in (Bu et al., 2022). For relatively simple cases (i.e. CIFAR-100 dataset), IF neuron requires approximately 4×4$\times$4 × time latency of parallel conversion to reach the comparable accuracy, which makes our method achieve 1523×15$\sim$ 23$\times$15 ∼ 23 × actual acceleration ratio.